determine the x and y intercept of the following line
3y-6=10x
THANK YOU GUYS SO MUCH

Answer by Cromlix(4381)

Hi there,
3y - 6 = 10x
3y = 10x + 6
y = 10/3x + 2
Cuts x axis when y = 0
0 = 10/3x + 2
10/3x = -2
10x = -6
x = -0.6 (-0.6, 0)x intercept
Cuts y axis when y = 0
y = 10/3x + 2
y = 10/3(0) + 2
y = 2 (0, 2)y intercept
Hope this helps :-)